OpenStudy (anonymous):

Rewrite as a logarithmic equation. e^(4)=y

OpenStudy (jdoe0001):

$$\large { a^x = \color{red}{y}\\ log_a\color{red}{y} =x } $$

OpenStudy (anonymous):

@jdoe0001 , thank you, but can you help me a little further on the steps?

OpenStudy (reemii):

what steps?

OpenStudy (anonymous):

@reemii for writing the e^4=y as a logarithmic I don't get the process he gave me.

OpenStudy (reemii):

the definition of the logarithm is \(\log_a x=y\) if and only if \(x=a^y\). if you see \(e^4=y\), you will apply a logarithm on both sides. there is a choice left to you, that is the choice of the "base" , the number \(a\) in the line above. Since you see \(e^4\), you choose \(a=e\), then you apply \(\log_e\) on both sides. (\(\log_ex=\ln x\))

OpenStudy (reemii):

LHS: \(\log_e (e^4) = 4\) (have you seen that?) RHS: \(\log_e y\) .. nothing else to say about it.

OpenStudy (reemii):

@_xcrvnts equation is \(4=\log_e y\). (=\(\ln y\))

OpenStudy (anonymous):